A countrys census iss tha population of tha country as 240 millian in 1300, 282 million in 200a, and 300 million in 2010. Fit a secand-dagras palynomial passing through thasc thrae points. (Lat the yaar 2000 bx- a and lat pix) represant tha populatian in millions.) pfx)- million Use this polynornial to predict the populatians in 2020 and in 2030 2020 2030million million
Answer #1

po282 lut ean 2000 スー-lo 10, 244) 33o a-lob CO o, 3u9 - 62ooa 2าะ loo (-0.03) +10b 2θ= lob (x 03 (20)+ 32o) In Lolo χ=20 (003